Find the value of a for which the eqn 


4 cosec 2 (π (a+x) ) + a2 - 4a = 0 has a real solution.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Sanchit

4 cosec 2 (π (a+x) ) + a2 - 4a = 0

4 cosec 2 (π (a+x) ) =- a2 + 4a 

L .H .S is always greater than 4 so for real solution

 -a2 + 4a ≥4

or (a-2)2≤0

so only a=2 is possible


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin